OpenStudy (anonymous):

Approximate the solution to 1 / square root of x^2 +1 = 1 / X + 5

OpenStudy (anonymous):

\[\LARGE \frac{1}{\sqrt{x^2+1}}=\frac{1}{x+5}\] use cross multiply and you'll get: \[\LARGE \sqrt{x^2+1}=x+5\] now square both sides... brb

OpenStudy (anonymous):

\[\LARGE (\sqrt{x^2+1}\;)^2=(x+5)^2 \] squaring them what do you get? :)

OpenStudy (anonymous):

\[X ^{2} + 1= (X+5)^{2}\] Then I would just solve for x right?

OpenStudy (anonymous):

hmm... yes but... \[\LARGE x^2+1=x^2+10x+25\] \[\LARGE \cancel{x^2}-\cancel {x^2}-10x=25-1\] \[\LARGE -10x=24\] \[\LARGE x=-\frac{24}{10 }\] simplify it.. :)

OpenStudy (anonymous):

Okay, thank you

OpenStudy (anonymous):

Glad to help :)